OpenStudy (anonymous):

(1+ cosx)raise to power 1/2 and limits are from pie/2 to 0

OpenStudy (anonymous):

can i one tell me

OpenStudy (anonymous):

are you looking for \[\int_{\frac{\pi}{2}}^0 \sqrt{1+\cos(x)}dx\] or \[\int_0^{\frac{\pi}{2}} \sqrt{1+\cos(x)}dx\]

OpenStudy (anonymous):

NO

OpenStudy (anonymous):

-2 *rad2* 1/rad2 -2

OpenStudy (anonymous):

then what on earth is it?

OpenStudy (anonymous):

i bet it is 2. just a guess though because this question certainly is eligible for the first karnak award of the day